OpenStudy (anonymous):

Simplify: (-4i+13)(-i-1) A. 17i-9 B. -9i-17 C. -17+9i D. 9+17i E. none of these. I'm getting E. none of these, the answer I'm coming out with is -17-9i

OpenStudy (michele_laino):

we have: \[(-4i+13)*(-i-1)=(-4i)*(-i-1)+13*(-i-1)=-4+4i-13i-13=-17-9i\]

OpenStudy (michele_laino):

oops: \[=-4+4i-13i-13=-17-9i\]

OpenStudy (michele_laino):

So B.

OpenStudy (vlery):

D

OpenStudy (vlery):

wait nvm B

OpenStudy (anonymous):

@Michele_Laino but the order of it is different. it's okay to switch it around like that?

OpenStudy (michele_laino):

yes, you can write a complex number in this form: a+ib, or in this form: ib+a, in both cases the complex number is the same

OpenStudy (michele_laino):

@clynnew